Find the equilibrium constant of the given reaction if the final concentration are as follows:

CH3COOH+CH3CH2OH<—>CH3CH2OOCCH3+H2O



CH3COOH= 0.302 M

CH3CH2OH= 0.428 M

CH3CH2OOCCH3= 0.665 M


1
Expert's answer
2021-06-14T13:35:37-0400

The equilibrium constant of the given reaction is-

K=[CH3CH2OOCCH3]/[CH3COOH] [CH3CH2OH]

Given,

CH3COOH= 0.302 M

CH3CH2OH= 0.428 M

CH3CH2OOCCH3= 0.665 M

Now,

Equilibrium constant(K)=(0.665)/(0.302)(0.428)

K=(0.665)/(0.129)=5.155

Hence,the equilibrium constant of the given reaction is 5.155